OpenStudy (ender):

I've spent all day trying to understand this and i really need help- (7*5*2 over 7*3)^2 * (5^0 over 2^-3)^3 * 2^-9 Can anyone help?

OpenStudy (rcp031):

U must have an Understanding of Order of Operations: Parenthesis (Grouping), Exponents Multiply and Divide (Left to Right) Add and Subtract (Left to Right)

OpenStudy (rcp031):

U if understand Order of Operations then u can solve such expression first step solve things under Parenthesis (Grouping) i.e, (7*5*2 over 7*3)^2=(10/3)^2=100/9 (5^0 over 2^-3)^3=(1/1/8)^3=8^3=512 second step Multiply (Left to Right) i.e, (100/9)*512*512 & do the math.

OpenStudy (needrehab):

Think About please excuse my dear aunt sally